# CliniBench: A Clinical Outcome Prediction Benchmark for Generative and Encoder-Based Language Models

### v1

@@ A.1 Prompts
We use the following prompt structure for each respective task setting:
@@
A.2 Dataset Preparation
We extract admission notes from MIMIC-IV-Note dataset Johnson et al. (2023b) and ICD discharge codes from MIMIC-IV dataset Johnson et al. (2023a) following van Aken et al. (2021).
We generate a stratified split of the datasets (s. Table˜1) into training, validation, and test sets, allocating 70% of the data for training, 10% for validation, and 20% for testing.
Additionally, we exclude labels that are not present across all three splits from the evaluation process.
For a more granular analysis of performance, we split labels into tertile bins based on the frequency of occurrence in the dataset: head (1st tertile), body (2nd tertile) and tail (3rd tertile).
A.3 JSON Generation Capabilities
Model #Parsable Avg. #Diag.
Qwen2 7B 0.00% -
Llama 3.1 8B 0.00% -
Mistral 7B 0.00% -
Mistral Nemo 0.00% -
Qwen2 72B 0.00% -
Llama 3.1 70B 0.00% -
Qwen2 7B Instruct 97.25% 19.91
Llama 3.1 8B Instruct 0.49% -
Mistral 7B Instruct 99.76% 18.16
Mistral Nemo Instruct 0.07% 20.00
Qwen2 72B Instruct 100.00% 20.12
Llama 3.1 70B Instruct 3.82% 20.16
A.4 Text-to-class Mapper Performance
Model Recall Precision MAP
all-mpnet-base-v2 9.99 6.02 10.94
S-PubMedBERT MSMarco 10.61 6.31 10.71
NeuML PubMedBERT 10.72 6.92 11.82
A.5 HPO and Threshold Tuning
[…]
Figure 6: Description of the threshold tuning function. Given a class $c$, we tune the thresholds using all observed logits ($T_{c}$) and all observed $F_{1}$ scores ($F_{c}$) on the validation set.
@@ A.5 HPO and Threshold Tuning
A.6 Additional Results
Model Data AUROC
BioMedBERT hosp ICD-9 90.12
ICD-10 88.01
icu ICD-9 83.59
ICD-10 80.38
GatorTronS hosp ICD-9 91.96
ICD-10 89.43
icu ICD-9 86.23
ICD-10 82.17
M2-BERT hosp ICD-9 85.82
ICD-10 81.05
icu ICD-9 77.99
ICD-10 75.08
Table 7: AUROC performance of the fine-tuned encoder models.
@@ A.6 Additional Results
Figure 7: Sankey diagram of adapted and ignored predictions for Qwen2-72B zero-shot and few-shot augmented experiments (5-shot BM25). Scores aggregated over all datasets.

### v2

@@ A.1 Prompts
For few-shot demonstrations, we provide the admission notes together with their annotated diagnosis short descriptions in the desired JSON format. We use the following prompt structure for each respective task setting:
@@
We hypothesize that providing the model with comprehensive information about the ICD ontology would enhance its performance. However, the full set of code descriptions exceeds the maximum context length of all evaluated models. Using the Qwen2.5 tokenizer, the descriptions of all possible ICD-9 codes require 167,679 tokens, and those of ICD-10 codes require 1,369,475 tokens. Consequently, we restrict the input to the task prompt, admission note and few-shot examples.
CoT implementation.
A.2 Construction of the Admission Notes
To avoid data leakage and maintain temporal consistency, we carefully select which sections of clinical notes to include based on whether the information would realistically be available at admission time.
Included sections with information available at admission time: Chief Complaint, Major Surgical or Invasive Procedure, Allergies, History of Present Illness, Past Medical History, Social History, Family History, Physical Exam at Admission, and Medication at Admission
Discarded sections that contain information not available at admission time: Physical Exam during Stay and at Discharge, Pertinent Results, Brief Hospital Course, Medication at Discharge, Discharge Disposition, Facility, Discharge Diagnosis, Discharge Condition, Discharge Instructions, and Followup Instructions
Sections that do not contribute to the task or prediction and are therefore discarded: Name, Unit No, Admission Date, Discharge Date, Date of Birth, Sex, Service, and Attending. This systematic approach ensures that our models are evaluated under realistic clinical conditions where only admission-time information is available for outcome prediction
A.3 HPO and Threshold Tuning
See Table˜4 for a breakdown of our hyperparameter optimization.
@@ A.3 HPO and Threshold Tuning
A.4 Retrieval Model Details
The retrieval evaluation enables us to gauge the impact of adding a generative LLM on top of the retrieved few-shot results. We evaluate the following semantic similarity retrievers based on the S-BERT architecture Reimers and Gurevych (2019): NeuML-PubmedBERT, jina-base-v2 günther-2023-jina, all-mpnet-base-v2444https://huggingface.co/sentence-transformers/all-mpnet-base-v2, all-mpnet-base-v2-neg Anschütz et al. (2023), BiomedBERT Gu et al. (2021), and S-PubMedBert-MS-MARCO Deka et al. (2022). For the latent outcome similarity retrieval, we use the CLS token from the last embedding layer of our fine-tuned baseline encoder (BioMedBERT) as a representation of the admission note.
@@ A.4 Retrieval Model Details
A.5 Computational Cost Analysis
ICU HOSP
Model #Parameters ICD 9 ICD 10 ICD 9 ICD 10
BioMedBERT 110M 1.17 0.73 5.23 2.89
GatorTronS 345M 3.45 2.16 15.41 8.53
M2 BERT 80M 1.12 0.72 4.97 2.80
Llama 3.1 8B 8B 103.08 64.68 454.58 252.1
Llama 3.1 70B 70B 901.93 565.97 3977.57 2206.30
Mistral 7B 7B 109.79 68.84 484.08 267.82
Mistral Nemo 13B 13B 181.99 113.66 802.34 441.28
Qwen 2 7B 7B 93.15 58.27 409.40 226.76
Qwen 2 72B 72B 599.37 958.10 2332.40 4210.96
Table 6: Inference cost in PFLOPs for each test split using the respective tokenizer of the model. We approximate the number of FLOPs based on Kaplan et al. (2020) by multiplying the number of tokens of each dataset with the number of parameters of the model multiplied by 2. Qwen2, due to its more efficient tokenizer, reduces compute costs compared to Llama 3.1 by almost 40%.
The computational costs differ dramatically between model types. Training a BERT-base model on the ICD-10 Hosp dataset requires about 9.51 PFLOPs per epoch Kaplan et al. (2020), or 28.53 PFLOPs when including the backward pass. With a maximum of 20 training epochs, this totals 570.6 PFLOPs for our experiments. By contrast, simply evaluating the ICD-10 Hosp test set with generative models is far more expensive: Qwen2-72B requires approximately 4210.96 PFLOPs, while Llama 3.1 8B needs 252.1 PFLOPs. The smaller BERT-based model only uses only 2.89 PFLOPs for inference on the same test set (see Table˜6). This demonstrates that generative models require orders of magnitude more computation than encoder-based models, even for inference on a small test set. These estimates ignore several real-world considerations, such as infrastructure requirements and hardware utilization. Larger generative models demand substantially more hardware during inference, which might remain underutilized between tasks, though these costs could be partially offset by leveraging techniques such as pruning and quantization.
Appendix B Additional Results
Model Data AUROC
BioMedBERT HOSP ICD-9 90.12
ICD-10 88.01
ICU ICD-9 83.59
ICD-10 80.38
GatorTronS HOSP ICD-9 91.96
ICD-10 89.43
ICU ICD-9 86.23
ICD-10 82.17
M2-BERT HOSP ICD-9 85.82
ICD-10 81.05
ICU ICD-9 77.99
ICD-10 75.08
Table 7: AUROC performance of the fine-tuned encoder models.
B.1 Text-to-class Mapper Performance.
We find that generative models often output diagnostic descriptions that exactly match the short or long descriptions of the respective ICD codes. This allows us to independently compare the performance of exact matched descriptions and codes that are mapped via the class mapper shown in Table˜9. For Qwen2-72B Instruct in the zero-shot scenario, only 5% of generated descriptions match ICD codes exactly; with 1, 3, and 5 few shots, this number rises to 32%, 51%, and 63% respectively. For the non-instruct fine-tuned version of the model, the number of exact matches rises even further to 56%, 70% and 76%. This indicates an adaption to the task through the provided few-shot examples. Furthermore, it shows that the model has seen the respective ICD short or long descriptions during pre-training and can correctly replicate them when it is provided with enough context. Among 5-shot predictions of Qwen2-72B Instruct, directly matched codes are correct 41% of the time, while mapped ones reach 20%. For Qwen2-7B, mapped codes retain 94% of the accuracy of directly matched ones (30% vs. 32%). While the accuracy of mapped codes is lower, their generated text contains enough relevant information that enables the class-mapper to resolve the correct code. This gap in performance is closely tied to the quality of the generated text. Accurate mappings result from meaningful clinical descriptions (e.g., Hypertensive heart disease, Hypotension), while nonsensical outputs (e.g., diagnosis 1, diagnosis 2, ___ etc.) almost never yield correct codes.
Model #Parsable Avg. #Diag.
Qwen2 7B 0.00% -
Llama 3.1 8B 0.00% -
Mistral 7B 0.00% -
Mistral Nemo 0.00% -
Qwen2 72B 0.00% -
Llama 3.1 70B 0.00% -
Qwen2 7B Instruct 97.25% 19.91
Llama 3.1 8B Instruct 0.49% -
Mistral 7B Instruct 99.76% 18.16
Mistral Nemo Instruct 0.07% 20.00
Qwen2 72B Instruct 100.00% 20.12
Llama 3.1 70B Instruct 3.82% 20.16
Model Recall Precision MAP
all-mpnet-base-v2 9.99 6.02 10.94
S-PubMedBERT MSMarco 10.61 6.31 10.71
NeuML PubMedBERT 10.72 6.92 11.82
B.2 Performance drop for longer inputs.
Figure 5: Micro $F_{1}$ over different sequence lengths averaged over all dataset splits. Encoder models have a reduced performance with sequences >850 tokens while the generative models tend have a more consistent performance on longer sequences. The token count varies across models due to differences in tokenizers.
B.3 ICU Results
Data Retriever #FS Rec. Prec. MAP MD Acc.
ICU Zero-Shot prompted 0 24.11 14.20 16.82 48.11
Random 1 17.67 13.95 16.06 40.80
3 15.90 13.27 16.00 41.22
5 14.24 13.45 15.77 40.66
BM25 1 19.97 15.29 17.46 47.27
3 18.89 15.89 19.09 51.18
5 17.42 15.73 20.48 52.06
PubMedBERTNeuML 1 18.25 13.83 17.16 45.59
3 17.13 13.97 18.43 50.50
5 16.10 14.48 19.92 52.34
BioMedBERTft 1 18.97 14.60 17.69 44.80
3 18.04 14.86 18.96 50.92
5 18.12 15.44 19.57 52.59
Gold Heuristic 1 21.31 17.79 19.75 48.45
3 22.72 20.24 26.88 59.15
5 22.83 21.21 29.33 61.66
Table 10: Few-shot performance of Qwen2-72B-Instruct averaged over ICD-9 and ICD-10 ICU splits using the depicted retrieval models. Metrics are macro averaged. BioMedBERTft is a custom fine-tuned encoder model on the respective training split of the dataset. (#FS: number of few-shot demonstrations, MD Acc.: Main diagnosis accuracy)
Figure 6: Macro recall and precision by model and classes grouped into tertiles by class frequency aggregated over both ICU splits (ICD-9 and ICD-10).
Zero-Shot prompted Zero-Shot prompted + CoT Few-Shot 5 Demonstrations
Data Model Rec. Prec. MAP MD Acc. F1 Rec. Prec. MAP MD Acc. F1 Rec. Prec. MAP MD Acc. F1
ICU Encoder-based Models
BioMedBERT-110M 19.60 17.99 32.44 72.83 22.00
BioMedBERT-110M-tuned 15.56 23.13 20.03 55.91 20.71
GatorTronS-345M 28.61 23.96 35.66 78.44 25.77
GatorTronS-345M-tuned 20.00 29.62 21.13 63.66 25.34
M2-80M 10.16 08.27 23.12 56.05 9.50
M2-80M-tuned 08.13 08.98 10.72 35.05 8.34
Generative LLM
Llama 3.1-8B 13.97 14.81 13.50 33.43 11.20 12.25 13.41 12.36 35.57 09.76 14.12 14.60 17.62 48.33 13.20
Llama 3.1-8B-Instruct 19.59 13.93 15.63 47.34 13.02 17.88 13.00 13.21 47.65 11.49 15.52 14.36 15.93 44.27 13.32
Llama 3.1-70B 15.40 12.44 12.30 33.60 10.48 18.73 11.87 12.98 40.81 11.22 13.74 12.10 16.54 50.85 12.13
Llama 3.1-70B-Instruct 20.31 13.36 14.70 50.37 12.55 21.24 12.66 15.85 50.18 12.08 17.06 15.38 16.24 44.90 14.47
Mistral-7B 08.84 13.49 09.94 25.66 08.08 00.34 02.63 04.58 01.38 00.12 11.08 11.92 14.40 43.16 10.79
Mistral-7B-Instruct 23.65 13.67 16.86 51.37 13.69 21.59 12.12 14.64 46.10 12.30 20.96 15.08 16.98 47.72 14.96
Mistral-Nemo-13B 13.49 14.31 13.55 31.63 10.86 11.11 11.63 10.86 37.51 08.23 11.31 12.73 15.13 44.95 11.00
Mistral-Nemo-13B-Instruct 19.64 11.97 13.87 48.39 11.58 19.03 11.96 12.46 48.54 10.78 11.24 13.35 15.24 43.28 11.37
Qwen2-7B 20.95 13.92 15.54 40.97 13.11 19.05 13.96 15.49 40.19 12.55 11.23 13.36 15.80 43.35 11.35
Qwen2-7B-Instruct 23.06 14.14 15.33 45.86 14.10 22.13 13.89 15.18 46.30 13.64 18.23 14.40 17.61 50.52 14.15
Qwen2-72B 22.55 14.06 15.89 43.28 13.70 20.32 13.14 15.57 40.93 12.56 15.27 13.80 17.39 49.40 13.31
Qwen2-72B-Instruct 24.11 14.20 16.82 48.11 14.41 22.78 14.10 16.47 47.16 14.12 17.42 15.74 18.75 52.06 14.77
Table 11: (MAP: Mean Average Precision, MD Acc.: Main Diagnosis Accuracy) Zero-shot performance on the ICU split over the top-20 predictions. The reported metrics are macro averaged. We aggregate ICD-9 and ICD-10 and report an unweighted mean. The dashed lines separate the encoder from the generative LLMs. CoT is not beneficial while instruct fine-tuning boosts performance. Threshold tuning decreases the performance of encoder models within the top-20 predictions. We depict the best performing models in bold and underline the best performing generative models in each category.